OpenStudy (anonymous):

Did I do it right? h(x)=3x-|4-x|. Find all solutions to h(a)=10+(5/2)a

OpenStudy (anonymous):

i think i do: h(a)=3(a)-|4-a| 10+(5/2)a=3a-|4-a|

OpenStudy (anonymous):

then would be: 10-(1/2)a=-(4-a) and, 10-(1/2)a=4-a on the right track?